# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ 7.4 Overall Evaluation (RQ1)
TabDDPM $0.550\scriptscriptstyle\pm\scriptstyle.001$ $0.230\scriptscriptstyle\pm\scriptstyle.002$ $0.720\scriptscriptstyle\pm\scriptstyle.003$ $0.031\scriptscriptstyle\pm\scriptstyle.001$ $0.130\scriptscriptstyle\pm\scriptstyle.002$ $0.064\scriptscriptstyle\pm\scriptstyle.003$ $0.543\scriptscriptstyle\pm\scriptstyle.001$ $0.164\scriptscriptstyle\pm\scriptstyle.002$ $0.141\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.030\scriptscriptstyle\pm\scriptstyle.001}$ $0.383\scriptscriptstyle\pm\scriptstyle.002$ $0.120\scriptscriptstyle\pm\scriptstyle.003$
GReaT $0.495\scriptscriptstyle\pm\scriptstyle.002$ $0.236\scriptscriptstyle\pm\scriptstyle.003$ $0.568\scriptscriptstyle\pm\scriptstyle.001$ $0.082\scriptscriptstyle\pm\scriptstyle.002$ $0.125\scriptscriptstyle\pm\scriptstyle.003$ $0.036\scriptscriptstyle\pm\scriptstyle.004$ $0.357\scriptscriptstyle\pm\scriptstyle.002$ $0.181\scriptscriptstyle\pm\scriptstyle.003$ $0.065\scriptscriptstyle\pm\scriptstyle.001$ $0.536\scriptscriptstyle\pm\scriptstyle.002$ $0.072\scriptscriptstyle\pm\scriptstyle.003$
@@
Fidelity Evaluation. We assess the data synthesis fidelity by applying the Wasserstein distance to both training $\mathcal{D}_{\text{train}}$ and test data $\mathcal{D}_{\text{test}}$. The results in Table 4 and Appendix Table D17 demonstrate that TabDDPM leads among HP synthesizers, while statistical methods (e.g., PGM) excel within the DP synthesizers category. To achieve differential privacy, all deep generative models experience a significant fidelity loss, whereas statistical methods maintain stable performance. This stability is further highlighted when comparing performance across training and test data, where statistical methods exhibit a lesser performance drop compared to deep generative models. This suggests that statistical methods are more robust in data synthesis. A closer examination of the fidelity of each type of attribute (detailed in Appendix Section D.1) reveals that statistical methods generally achieve better fidelity performance on categorical attributes but fall short with numerical ones. This discrepancy can be attributed to the inherent design of statistical methods for discrete spaces, making them less effective for numerical data. Despite DDPM being enhanced with the multinomial diffusion process to better model discrete distributions, it still does not match the performance of statistical methods in terms of categorical fidelity.
@@
Privacy Evaluation. Table 5 presents the privacy assessment for HP synthesizers. Contrary to the fidelity evaluation findings, CTGAN, which exhibited the lowest fidelity performance, provides superior privacy protections against membership disclosure. Statistical methods, such as PGM, also demonstrate commendable empirical privacy protections. Nonetheless, the unsatisfied results of strong synthesis algorithms like TabDDPM and GReaT reveal their vulnerability to membership disclosure.
@@
Utility Evaluation. The utility of data synthesis is assessed by performing downstream tasks on the synthetic datasets and measuring with the proposed metrics (machine learning affinity and query errors), as shown in Table D28 and Table D29 (in Appendix). In the context of machine learning tasks, TabDDPM excels among HP synthesizers, which contributes to its class-conditional framework that learns label dependencies during its training process. However, this advantage diminishes when adding random noise to ensure privacy, where PGM takes the lead with its stable and superior performance. The outcomes for range (point) query tasks echo the fidelity evaluation results, where TabDDPM leads in scenarios devoid of provable privacy guarantees, and statistical methods can surpass other methods under DP constraints.
@@
Figure 6: Impact of privacy budget $\epsilon$ on Bean dataset. “Train Fidelity” means the fidelity is evaluated on train dataset $\mathcal{D}_{\text{train}}$, and “Test Fidelity” denotes evaluated on test dataset $\mathcal{D}_{\text{test}}$. “MLA” is the proposed machine learning affinity.
7.5 In-depth Analysis (RQ2)
@@
In this section, we delve into the underlying reasons for the observed performances. Specifically, we employ the proposed fidelity metrics as tools to auditing the learning process of different synthesizers and explore the impact of differing privacy budgets on different DP synthesizers.
@@
Figure 7: Auditing the learning process of CTGAN and TabDDPM with fidelity metrics on the Bean dataset. The fidelity of numerical-numerical marginals is omitted since the Bean dataset only has one numerical attribute.
@@
Why Dose CTGAN Perform Poorly? CTGAN is widely recognized as a strong synthesis algorithm in many studies [32, 7, 71]. Despite its popularity, our evaluation reveals that it produces the lowest quality synthetic data. This discrepancy prompts an investigation into the reasons behind CTGAN’s apparent underachievement.
@@
To address this, we scrutinize CTGAN’s learning trajectory, particularly evaluating the fidelity across different marginal types (categorical, numerical, and mixed) throughout training. Our findings illustrated in Figure 7(a), reveal an unexpected stagnation in the improvement of Wasserstein distances for both numerical and categorical marginals throughout the entire training period. As such, the quality of synthetic data completely relies on data preprocessing, which employs a variational Gaussian mixture model to approximate numerical data and conditional sampling for categorical attributes. The closer the data distribution to Gaussianity, the better the fidelity. Yet, most tabular data display complexities beyond simple Gaussian distributions [22], leading to CTGAN’s subpar performance. Furthermore, this phenomenon sheds light on CTGAN’s strong empirical privacy protections. The nearly random (and unlearnable) nature of the output does not depend on any specific training data presence. This aspect, while beneficial for privacy, underscores the challenges in generating high-fidelity synthetic data using CTGAN.
@@
Why Dose TabDDPM Excel? A notable finding from our evaluations is TabDDPM’s capability to maintain high fidelity of low-dimensional marginals. This contradicts with previous statements that deep generative models generally fail for tabular data synthesis [40, 58]. To better illustrate this, we utilized fidelity metrics to audit TabDDPM’s learning process, as depicted in Figure 7(b). A rapid decrease of the Wasserstein distance for all types of marginals is observed during the training, indicating the ability of TabDDPM to model both numerical and categorical attributes. We contribute the efficacy of TabDDPM to the model architecture and categorical distribution modeling: Diffusion models are theoretically proved to minimize the Wasserstein distance between synthetic and real datasets [33]. This represents a methodological superiority compared to other generative models, which primarily aim at reducing the Kullback–Leibler divergence. Moreover, the integration of the multinomial diffusion process addresses the challenge of modeling discrete spaces for conventional diffusion models. However, TabDDPM faces serious privacy disclosure risks which are ignored by previous work [32], and directly applying differential privacy (i.e., TableDiffusion) would greatly impact the model’s ability to generate authentic data. Despite these challenges, we deem that diffusion-based methods represent a promising frontier for data synthesis.
@@
Large Language Models Are Semantic-aware Synthesizers. We also notice that the recently emerged LLM-based synthesizer (i.e., GReaT) also shows competitive performance for data synthesis, which is capable of modeling semantic knowledge within attributes. This capability enables GReaT to excel in datasets characterized by explicit semantics. A good example is the Adult dataset, which comprises detailed personal information (e.g., age and income), and GReaT achieves the best performance on machine learning predictions among all HP synthesizers. Nevertheless, this advantage diminishes in cases where the attributes’ semantics are unclear. This is evident in the Phishing dataset, where all attributes’ values are preprocessed into integers and GReaT’s performance is notably weaker. Given the rapid development of LLM and the inherent rich semantics of most tabular data, it may become a new paradigm for realistic data synthesis.
@@
The Impact of Privacy Budget. We conduct experiments on different privacy budgets to further analyze the impact of differential privacy for data synthesis. Specifically, we vary $\epsilon$ from $0.5$ to $16$, and evaluate the fidelity and utility of synthetic data, as shown in Figure 6. It was observed that statistical methods (e.g., PGM) demonstrate robust performance even with the small privacy budget ($\epsilon=0.5$), whereas deep generative models generally require a much larger privacy budget (e.g., 8) to yield similar outcomes. This is reasonable since statistical methods only rely on the estimation of a small set of marginals. In contrast, deep generative models aim to directly capture the entire joint distribution, making them inherently more susceptible to random perturbations.
@@
Table 6: Average performance improvements (%) when tuning with our framework.
Models Fidelity $\uparrow$ Utility normal-$\uparrow$
@@
7.6 Effectiveness of proposed Metrics (RQ3)
In this section, we investigate the efficacy of proposed metrics by conducting a series of comparison studies.
Impact of Model Tuning Phase. We initiated our investigation by examining the impact of the proposed tuning objective in Section 6.1. The data presented in Table 6 highlights the average performance improvements across all datasets. Generally, the tuning objective can consistently improve both the fidelity and utility of synthetic data for all synthesis algorithms. We also notice that the improvement in utility surpasses those in fidelity for statistical methods, while deep generative models exhibit the reverse trend. In addition, the tuning phrase is exceptionally beneficial for TabDDPM, yielding significant improvements in both fidelity and utility. Conversely, GReaT and TVAE show a greater resilience to model tuning.
Analysis of Membership Disclosure Score. We investigate the proposed membership disclosure score (MDS) by employing it to evaluate the empirical privacy protections of DP synthesizers. Specifically, we set the privacy budget $\epsilon$ from 1 to infinity and calculate the corresponding MDS. The results on the Adult and Phishing dataset are shown in Figure 8. It is observed that MDS can indicate the change in privacy budget, and a larger privacy budget would lead to higher membership disclosure risks (thus larger MDS). Interestingly, we find that statistical methods can provide reasonable empirical privacy protections, even without DP guarantees (i.e., $\epsilon=\infty$). On the contrary, deep generative models like TableDiffusion may need a smaller privacy budget to reduce empirical privacy risks. This may be because statistical methods are inherently more private since they only rely on statistics for synthesis. We leave it as an open question for future research.
8 Key Takeaways
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. Despite the plethora of synthesis algorithms available, a comprehensive understanding of their strengths and weaknesses remains elusive due to the lack of principled evaluation metrics. In this paper, we examine and critique existing evaluation metrics, and introduce a principled and systematic framework for assessing data synthesis, as well as a new suite of evaluation criteria in terms of fidelity, privacy, and utility. We also provide a unified tuning objective that can consistently improve the performance of synthesis models and ensure that evaluations are grounded.
In the following, we summarize important insights obtained from our evaluations and provide some actionable advice to future data synthesis practitioners:
@@ 8 Key Takeaways
Figure 8: Empirical privacy evaluation with membership disclosure score (MDS) on DP synthesizers.
  • No Free lunch in data synthesis. Currently, the performance gap between HP and DP synthesizers is quite large. No synthesizer can achieve the best performance while maintaining good privacy protections (even empirically).
  • Model tuning is indispensable for synthesizers. we find that the performance of all synthesizers can be boosted with our proposed tuning objective compared to the default settings. Therefore, tuning is crucial to reach the right conclusions.
  • Diffusion models are strong tabular synthesizers. We observe that diffusion models (i.e., TabDDPM) are capable of synthesizing highly authentic tabular data, significantly outperforming other synthesis algorithms. However, the privacy issue of diffusion models needs further investigation.
  • Statistical methods are robust (private) synthesizers. While many studies consider statistical methods to be weak to compare [68, 32, 34, 31], we observe that they achieve good and stable performance under DP constraints. We also find that statistical methods may be inherently more private than deep generative methods, which leaves an open question for further exploration.
  • Large language models are promising synthesizers. GReaT is the first LLM-based tabular synthesis, which utilizes fine-tuning of a pre-trained LLM on textually encoded tabular data. Despite the simplicity, it shows impressive performance on datasets that consist of rich semantic attributes.
  • CTGAN is too weak to be a practical synthesizer. Our analysis highlights that CTGAN falls short as an effective synthesizer, with no discernible improvement in synthetic data quality observed during its training phase. Therefore, a synthesizer’s performance should not be judged solely in comparison to CTGAN.
Our principled assessment leads us to conclude that recently emerged generative models achieve impressive performance for tabular data synthesis and offer new directions in this field. At the same time, several critical challenges are also revealed (e.g., privacy issues of diffusion models, gaps between DP and HP synthesizers, etc.). Our evaluation metrics and framework play a crucial role in elucidating the advancements in data synthesis as well as serve as a benchmark for future works.

### v2

@@ 7.2. Experimental Setups
TabDDPM $0.550\scriptscriptstyle\pm\scriptstyle.001$ $0.230\scriptscriptstyle\pm\scriptstyle.002$ $0.702\scriptscriptstyle\pm\scriptstyle.003$ $0.031\scriptscriptstyle\pm\scriptstyle.001$ $0.130\scriptscriptstyle\pm\scriptstyle.002$ $0.064\scriptscriptstyle\pm\scriptstyle.003$ $0.543\scriptscriptstyle\pm\scriptstyle.001$ $0.164\scriptscriptstyle\pm\scriptstyle.002$ $0.141\scriptscriptstyle\pm\scriptstyle.003$ $\mathbf{0.030\scriptscriptstyle\pm\scriptstyle.001}$ $0.383\scriptscriptstyle\pm\scriptstyle.002$ $0.120\scriptscriptstyle\pm\scriptstyle.003$
GReaT $0.495\scriptscriptstyle\pm\scriptstyle.002$ $0.236\scriptscriptstyle\pm\scriptstyle.003$ $0.568\scriptscriptstyle\pm\scriptstyle.001$ $0.082\scriptscriptstyle\pm\scriptstyle.002$ $0.125\scriptscriptstyle\pm\scriptstyle.003$ $0.036\scriptscriptstyle\pm\scriptstyle.004$ $0.357\scriptscriptstyle\pm\scriptstyle.002$ $0.181\scriptscriptstyle\pm\scriptstyle.003$ $0.065\scriptscriptstyle\pm\scriptstyle.001$ $0.534\scriptscriptstyle\pm\scriptstyle.002$ $0.072\scriptscriptstyle\pm\scriptstyle.003$
SELF $0.732\scriptscriptstyle\pm\scriptstyle.000$ $0.498\scriptscriptstyle\pm\scriptstyle.000$ $0.718\scriptscriptstyle\pm\scriptstyle.000$ $0.178\scriptscriptstyle\pm\scriptstyle.000$ $0.426\scriptscriptstyle\pm\scriptstyle.000$ $0.257\scriptscriptstyle\pm\scriptstyle.000$ $0.659\scriptscriptstyle\pm\scriptstyle.000$ $0.318\scriptscriptstyle\pm\scriptstyle.000$ $0.284\scriptscriptstyle\pm\scriptstyle.000$ $0.267\scriptscriptstyle\pm\scriptstyle.000$ $0.537\scriptscriptstyle\pm\scriptstyle.000$ $0.342\scriptscriptstyle\pm\scriptstyle.000$
Baselines. We introduce additional baselines to better understand the performance of synthesizers on the proposed evaluation metrics. For fidelity, the HALF baseline randomly divides the real data into two equal parts: one serving as the training dataset $\mathcal{D}$, and the other as the synthetic data $\mathcal{S}$. Since the two are drawn from the same distribution, this serves as an upper bound on the fidelity scores. For HISTOGRAM, the synthetic dataset is generated by PGM with only one-way marginals (i.e., histograms). Thus in the synthetic dataset, all attributes are independent from each other. We use HISTOGRAM as the empirical lower bounds of fidelity.
For privacy, we introduce the SELF baseline, which uses a direct copy of the real data as the synthetic data, establishing the worst privacy protection (lower bound). By the definition of MDS, the perfect privacy-preserving synthesizer would achieve a score of 0, which is the upper bound of privacy.
Having these lower and upper bounds for fidelity and privacy enables us to better interpret the scores of different synthesizers. We do not include baselines for utility metrics (i.e., machine learning affinity and query errors), since their scores are ratios that can naturally reflect the relative deviation from the ground truth.
7.3. Overall Evaluation (RQ1)
Overview. Figure 2 and Figure 4 report the overview ranking results for HP and DP synthesizers, respectively. For HP synthesizers, TabDDPM exhibits superior fidelity and utility, albeit at the expense of compromising privacy. Statistical methods like PrivSyn also achieve impressive fidelity, surpassing complex deep generative models like TVAE and GReaT. Conversely, CTGAN, the most popular HP synthesizer, shows the least satisfactory results in synthetic data quality but offers better privacy protection.
In terms of DP synthesizers, statistical methods remain effective in both fidelity and utility. The performance of deep generative models drops significantly to ensure differential privacy. Even the strongest model (i.e., TableDiffusion) underperforms statistical approaches by a large margin, which starkly contrasts with its performance in the HP context, indicating a pronounced impact of privacy constraints on deep generative models.
We also visualize the synthetic and real data by using the t-SNE technique (Van der Maaten and Hinton, 2008), as depicted in Figure 3 and Figure 5. The visualizations generally reveal a considerable overlap between synthetic and real data points, demonstrating the synthesizers’ ability to generate data resembling the original. One interesting observation is that PrivSyn is the only method that does not have clear cluster patterns. This is because PrivSyn is the only non-parametric approach that directly modifies the synthetic records to match the marginal distribution, potentially resulting in less realistic records.
Table 5. Machine learning affinity of data synthesis. The lower value means the performance of synthetic data is more similar to real data. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The best result of each category is in bold.
[…]
Table 6. Query error of data synthesis. A lower value means a smaller query error. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The best result of each category is in bold.
@@ 7.3. Overall Evaluation (RQ1)
Fidelity Evaluation. We assess the data synthesis fidelity by applying the Wasserstein distance to both training $\mathcal{D}_{\text{train}}$ (shown in Table 3) and test data $\mathcal{D}_{\text{test}}$ (included in technical report version due to similar results). The results demonstrate that TabDDPM leads among HP synthesizers (nearly reaches the empirical upper bound), while statistical methods (e.g., PGM) excel within the DP synthesizers category. To achieve differential privacy, all deep generative models experience a significant fidelity loss, whereas statistical methods maintain stable performance. This stability is further highlighted when comparing performance across training and test data, where statistical methods exhibit a lesser performance drop compared to deep generative models. A closer examination of the fidelity of each type of attribute (detailed in technical report version) reveals that statistical methods generally achieve better fidelity performance on categorical attributes but fall short with numerical ones. This discrepancy can be attributed to the inherent design of statistical methods for discrete spaces, making them less effective for modeling numerical attributes.
@@
Privacy Evaluation. Table 4 presents the privacy assessment for HP synthesizers. Contrary to the fidelity evaluation findings, CTGAN, which exhibited the lowest fidelity performance, provides the best privacy protections against membership disclosure. Statistical methods, such as PGM, also demonstrate commendable empirical privacy protections. Nonetheless, the unsatisfied results of strong synthesis algorithms like TabDDPM and GReaT reveal their vulnerability to membership disclosure.
@@
Utility Evaluation. The utility of data synthesis is assessed by performing downstream tasks on the synthetic datasets and measuring with the proposed metrics (i.e., machine learning affinity and query errors), as shown in Table 5 and Table 6. In the context of machine learning tasks, TabDDPM excels among HP synthesizers, which contributes to its class-conditional framework that learns label dependencies during its training process. However, this advantage diminishes when adding random noise to ensure privacy, where PGM takes the lead with its robust and superior performance. The outcomes for range (point) query tasks echo the results of fidelity evaluation, where TabDDPM shows superior performance in HP settings, and statistical methods (i.e., PGM) can surpass other methods under DP constraints.
@@
Figure 6. Impact of privacy budget $\epsilon$ on Bean dataset. The lower score indicates higher fidelity/utility. “Train (test) Wasserstein Distance” denotes the fidelity is evaluated on train (test) dataset $\mathcal{D}_{\text{train}}$ ($\mathcal{D}_{\text{test}}$). “MLA” is the proposed utility metric.
7.4. In-depth Analysis (RQ2)
@@
In this section, we delve into the underlying reasons for the observed performances. Specifically, we employ the proposed fidelity metrics as tools for auditing the synthesizers’ learning process and explore the impact of differing privacy budgets on DP synthesizers.
@@
Figure 7. Auditing the learning process of CTGAN and TabDDPM with fidelity metrics on the Bean dataset. The fidelity of numerical-numerical marginals is omitted since the Bean dataset only has one numerical attribute.
@@
Why Does CTGAN Perform Poorly? CTGAN is widely recognized as a strong synthesizer in many studies (Kotelnikov et al., 2023; Borisov et al., 2023; Zhao et al., 2021). Despite its popularity, our evaluation reveals that it produces the lowest quality synthetic data. This discrepancy prompts an investigation into the reasons behind CTGAN’s apparent underachievement.
@@
To address this, we scrutinize CTGAN’s learning trajectory, particularly evaluating the fidelity across different marginal types throughout training. Our findings illustrated in Figure 7(a), reveal an unexpected stagnation in the improvement of Wasserstein distances for both numerical and categorical marginals throughout the entire training period. As such, the quality of synthetic data completely relies on data preprocessing, which employs a variational Gaussian mixture model to approximate numerical data and conditional sampling for categorical attributes. The closer the data distribution to Gaussianity, the better the fidelity. Yet, most tabular data display complexities beyond simple Gaussian distributions (Gorishniy et al., 2021), leading to CTGAN’s subpar performance. Furthermore, this phenomenon sheds light on CTGAN’s strong empirical privacy protections. The nearly random (and unlearnable) nature of the output does not depend on the presence of any individual training data. This aspect, while beneficial for privacy, underscores the challenges in generating high-fidelity synthetic data using CTGAN.
@@
Why Does TabDDPM Excel? A notable finding from our evaluations is TabDDPM’s capability to achieve high fidelity. This contradicts with previous statements that deep generative models generally fail for tabular data synthesis (McKenna et al., 2022; Tao et al., 2021). To better illustrate this, we utilized fidelity metrics to audit TabDDPM’s learning process. From the results depicted in Figure 7(b) we observe a rapid decrease in the Wasserstein distance for all types of marginals, indicating the ability to model both numerical and categorical attributes. We contribute its success to the design of the model architecture. Diffusion models are proved to minimize the Wasserstein distance between synthetic and real data (Kwon et al., 2022). This represents a methodological superiority compared to other generative models, which primarily aim at reducing the Kullback–Leibler divergence. Moreover, integrating the multinomial diffusion process addresses the challenge of modeling discrete spaces for conventional diffusion models. However, TabDDPM faces serious privacy disclosure risks which are ignored by previous work (Kotelnikov et al., 2023), and directly applying DP (i.e., TableDiffusion) would greatly impact the quality of synthetic data. Despite these challenges, we deem that diffusion-based methods represent a promising frontier for data synthesis.
@@
Large Language Models Are Semantic-aware Synthesizers. We also notice that the recently emerged LLM-based synthesizer (i.e., GReaT) also shows competitive performance, especially on datasets that consist of rich semantic attributes. For example, the Adult dataset comprises detailed personal information (e.g., age, income, and relationship), and GReaT achieves the best performance on machine learning predictions among HP synthesizers. Nevertheless, this advantage diminishes in cases where the attributes’ semantics are unclear. This is evident in the Phishing dataset, where all attributes’ values have been preprocessed into integers and GReaT’s performance is notably weaker. Given the rapid development of LLM and the inherent rich semantics of most tabular data, it may become a new paradigm for realistic data synthesis.
(a) Privacy Metric Comparison
(b) Machine Learning Performance
Figure 8. (a) Comparing different privacy metrics on Adult dataset. Existing metrics are uninformative and only MDS can accurately detect the privacy risks of different synthesizers. (b) Performance behaviors of HP synthesizers on Magic dataset. “LR” denotes Linear Regression, “DT” is Decision Tree, and “RF” means Random Forest.
@@
The Impact of Privacy Budget. We conduct experiments on different privacy budgets to further analyze the impact of differential privacy for data synthesis. Specifically, we vary $\epsilon$ from $0.5$ to $16$, and evaluate the fidelity and utility of synthetic data, as shown in Figure 6. It is observed that statistical methods (e.g., PGM) demonstrate robust performance even with the small privacy budget (e.g., $\epsilon=0.5$), whereas deep generative models generally require a much larger privacy budget (e.g., $\epsilon=8$) to yield similar outcomes. This is reasonable since statistical methods only rely on the estimation of a small set of marginals. In contrast, deep generative models aim to directly capture the entire joint distribution, making them inherently more susceptible to random perturbations.
7.5. Effectiveness of proposed Metrics (RQ3)
Effectiveness of Membership Disclosure Score (MDS). Figure 8(a) shows the performance of different privacy metrics (i.e., DCR (Zhao et al., 2021), MI attack (Stadler et al., 2022) and MDS, detailed implementations are included in technical report version). To evaluate the privacy protection, we add a DP synthesizer (i.e., TableDiffusion with $\epsilon=1$) for reference. A good privacy metric should differentiate the privacy risks between TableDiffusion and its HP equivalent (i.e., TabDDPM). However, the results of DCR indicate all synthesizers have the same level of privacy protection, and the attack accuracy of MI attacks is rather low for all synthesizers. On the contrary, MDS can accurately detect the privacy risks of different synthesizers. For example, it indicates the membership disclosure risks of TabDDPM and GReaT are remarkably high, and applying DP to the synthesizer (i.e., TableDiffusion) can largely alleviate the disclosure risks.
Effectiveness of Machine Learning Affinity. The performance of various data synthesizers fluctuates significantly across different machine learning models, as illustrated in Figure 8(b). Such variations in performance underscore the impact of the choice of evaluation models. For instance, while PrivSyn is ranked third when evaluated using linear regression, it falls to fifth when assessed with decision trees. The proposed MLA addresses this challenge by quantifying relative discrepancies in performance trends, thereby facilitating a more nuanced and consistent utility evaluation.
@@ 7.5. Effectiveness of proposed Metrics (RQ3)
Table 7. Average performance improvements (%) on fidelity and utility when tuning with the proposed tuning objective.
Models Fidelity $\uparrow$ Utility $\uparrow$
@@
Impact of Model Tuning Phase. We showcase the average performance improvements by incorporating the tuning phase in Section 6.1. Generally, the tuning objective can consistently improve synthetic data’s fidelity and utility for all synthesizers. For statistical methods, the improvement in utility surpasses those in fidelity, while deep generative models exhibit the reverse trend. In addition, the tuning phrase is exceptionally beneficial for TabDDPM, yielding significant improvements in both fidelity and utility.
Empirical Privacy Analysis with MDS. We further investigate the empirical privacy protections of DP synthesizers by applying the proposed privacy metric (i.e., MDS). Specifically, we set the privacy budget $\epsilon$ from 1 to infinity and calculate the corresponding MDS. The results on the Adult and Phishing dataset are shown in Figure 9. It is observed that a larger privacy budget would lead to higher membership disclosure risks (thus larger MDS). Interestingly, we find that statistical methods can provide reasonable empirical privacy protections, even without DP guarantees (i.e., $\epsilon=\infty$). On the contrary, deep generative models like TableDiffusion need a smaller privacy budget for good empirical privacy protection. This may be because statistical methods are inherently more private since they only rely on statistics for synthesis. We leave it as an open question for future research.
@@
Figure 9. Empirical privacy evaluation with membership disclosure score (MDS) on DP synthesizers.
8. Discussion and Key Takeaways
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. Despite the plethora of available data synthesizers, a comprehensive understanding of their strengths and weaknesses remains elusive, due to the limitations of existing evaluation metrics. In this paper, we examine and critique existing metrics, and introduce a systematic framework as well as a new suite of evaluation criteria for assessing data synthesizers. We also provide a unified tuning objective that can consistently improve the performance of synthesizers and ensure that evaluation results are less affected by accidental choices of hyperparameters.
Our results identify several guidelines for practitioners seeking the best performance for their task and facing the daunting task of selecting and configuring appropriate synthesizers.
  • Model tuning is indispensable. Tuning hyperparameters can significantly improve synthetic data quality, especially for deep generative models.
  • Statistical methods should be preferred for applications where privacy is paramount. PRM and PrivSyn achieve the best fidelity among DP synthesizers, and they also offer good empirical privacy protection even in HP settings.
  • Diffusion models provide the best fidelity and utility. Practitioners are suggested to use diffusion models (e.g., TabDDPM) for tabular synthesis when the quality of synthetic data is the priority over privacy due to their impressive ability to generate highly authentic synthetic data.
  • Deep generative models can be tailored for specific tasks. The flexible design spaces of deep generative models make them suitable for scenarios where the applications of the synthetic data are known in advance (e.g., machine learning prediction). In addition, the LLM-based synthesizer, GReaT, is particularly effective at preserving semantic information in synthetic data.
Our systematic assessment leads us to conclude that recently emerged generative models achieve impressive performance on tabular data synthesis and offer new directions in this field. At the same time, several critical challenges are also revealed (e.g., privacy issues of diffusion models, performance gaps between DP and HP synthesizers, etc.). Our evaluation metrics and framework play a crucial role in elucidating the advancements in data synthesis as well as serve as a benchmark for future works.